Repeat 10.00-mL aliquots of an approximately 0.05 M hydrochloric acid solution were titrated with a 0.03484 M NaOH solution. The mean titration volume of the NaOH solution for an acceptable set of titrations was 9.60 mL. Calculate the exact molarity of the hydrochloric acid solution.
Answer #1

HCl + NaOH ----> NaCl + H2O

HCl                                     NaOH

M1 = ?                                M2 = 0.03484 M

V1 = 10.0 mL                       V2 = 9.60 mL

M1* 10.0 = 0.03484*9.60

M1 = (0.03484*9.6)/10

      = 0.0334464 M

      = 0.03345 M

exact molarity of the hydrochloric acid solution = 0.03345 M